OpenStudy (shelld93):

Rearrage to make y the subject. I need working as well please ! My online class never went into so much detail ! k = -3y + 4.3 - 6y

OpenStudy (imstuck):

First you have to isolate the -6y term by adding 3y to both sides ad then subtracting 4.3 from both sides. This leaves us with k + 3y - 4.3 = -6y. Now divide everything by -6 to get k + 3y - 4.3/ -6 = y. Since the 6 is negative, and I believe I am correct in saying this, the signs of the terms in the numerator will change. The answer, then, would be -k - 3y + 4.3/6 = y.

OpenStudy (shelld93):

so k = -3y + 4.3 - 6y k + 3y = -3y + 4.3 - 6y + 3y k + 3y - 4.3 = 4.3 - 6y - 4.3 k + 3y - 4.3/-6 = -6y/-6 -k - 3y + 4.3/-6 = y ?

OpenStudy (shelld93):

sorry last line is positive 6

OpenStudy (perl):

still need help?

OpenStudy (perl):

it is easier if you first collect like terms on the right side

OpenStudy (shelld93):

Sorry i wasnt here ! yeah, based on what the other person told me i ended up with k = -3y + 4.3 - 6y k + 3y = -3y + 4.3 - 6y + 3y k + 3y - 4.3 = 4.3 - 6y - 4.3 k + 3y - 4.3/-6 = -6y/-6 -k - 3y + 4.3/6 = y i that right? do brackets need to go anywhere ?

OpenStudy (perl):

i would prefer to start over, is that ok ?

OpenStudy (shelld93):

yeah of course

OpenStudy (shelld93):

thank you for helping me :)

OpenStudy (perl):

on the right side we can collect like terms

OpenStudy (perl):

k = -3y + 4.3 - 6y k = -9y + 4.3 ok so far ?

OpenStudy (shelld93):

yep

OpenStudy (perl):

now subtract 4.3 from both sides

OpenStudy (perl):

k = -3y + 4.3 - 6y k = -9y + 4.3 k - 4.3 = -9y

OpenStudy (perl):

make sense so far?

OpenStudy (shelld93):

yeah im following well :)

OpenStudy (perl):

now divide both sides by -9 ,

OpenStudy (perl):

to isolate the y

OpenStudy (perl):

k = -3y + 4.3 - 6y k = -9y + 4.3 k - 4.3 = -9y (k - 4.3) / -9 = y

OpenStudy (shelld93):

so much easier thank you so much

OpenStudy (shelld93):

I have another if you wouldn't mind ?

OpenStudy (perl):

ok

OpenStudy (shelld93):

t = -8(y-1)^2

OpenStudy (shelld93):

can i have a go first and you tell me where i mess up if thats okay ?

OpenStudy (shelld93):

actually i dont like that idea >.< im lost already

OpenStudy (perl):

so what was the original question

OpenStudy (shelld93):

t = -8(y-1)^2

OpenStudy (perl):